# One Look is Enough: A Novel Seamless Patchwise Refinement for Zero-Shot Monocular Depth Estimation Models on High-Resolution Images

### v2

@@ Abstract
Figure 1: Qualitative comparison of patch-based depth estimation models (PatchFusion [23], PatchRefiner [24] and PRO (Ours)) for high-resolution images. PatchRefiner [24] and PRO (Ours) also visualize the residuals to effectively highlight depth discontinuity issues in the 4th and 6th columns. All models perform depth estimations on $4\times 4$ patches of input images. Zoom-in can help distinguish the grid lines (depth discontinuities) of the $4\times 4$ patches. Our proposed PRO achieves smooth boundary transitions without depth discontinuity artifacts along the grid boundaries.
@@ 1 Introduction
To address the aforementioned problems, we propose Patch Refine Once (PRO), a novel refinement model that enables DE models to generate accurate depth predictions with only a single refinement per patch at inference time (i.e.,“One Look”). Our PRO consists of two strategies: Grouped Patch Consistency Training and Bias Free Masking. Fig. 1 shows a qualitative comparison of patch-based depth estimation models for high-resolution images. As shown, our PRO yields depth refinement results with minimal depth discontinuities along grid boundaries, whereas two recent state-of-the-art (SOTA) models still exhibit noticeable boundary artifacts. Fig. 2 illustrates a comparison of performance and efficiency between recent SOTA depth refinement models and our PRO. As observed, our PRO provides the best performance in terms of $\mathrm{D}^{3}\mathrm{R}$, AbsRel and inference speed. Our contribution is summarized as:
@@ 4.2 Implementation Details
We adopt the pretrained DepthAnythingV2 (DA2) [47] as the baseline model ($\Psi$ in Fig. 4). PRO, PatchFusion [23], and PatchRefiner [24] are (re)trained based on DA2, while BoostingDepth [26] uses DA2 without additional training. The input resolution to $\Psi$ is fixed at $518\times 518$. When we use the GPCT strategy, adjacent patches are cropped with an overlap of 224 pixels, ensuring an overlap of $h\times 224$ for horizontally adjacent patches and $224\times w$ for vertically adjacent patches, where $h$ and $w$ denote the height and width of the patch, respectively. In our BFM, the dilation kernel size is empirically set to (10, 20). Training is conducted on 7,592 UnrealStereo4K samples for 8 epochs with a batch size of 64 (with gradient accumulation), taking 10 hours on a single RTX 4090 GPU. Additional details are described in supplementary material.
@@ 4.3 Performance Comparison
At test time, each input image is divided into a $4\times 4$ grid of patches. The depth of each patch is refined individually, and the refined patches are then reassembled to generate the final depth map.
@@ Appendix A Extra Experiments
$3\times 3$ 0.0305 0.994 0.0423 0.985 0.0288 0.995 0.0805 0.125
$4\times 4$ 0.0304 0.994 0.0422 0.985 0.0287 0.996 0.0803 0.156
$5\times 5$ 0.0305 0.994 0.0423 0.985 0.0288 0.996 0.0797 0.162
@@ Appendix B Architecture of the Fusion Module
The encoder takes $\mathbf{P}^{i}$, $\mathbf{D}_{\mathrm{f}}^{i}$, and $\mathsf{ROI}(\mathbf{D}_{\mathrm{c}},\mathbf{P}^{i})$ as inputs and produces a set of five intermediate features, denoted as $\mathbf{F}_{\mathrm{enc}}=\{\mathbf{f}_{\mathrm{enc},j}^{i}\}_{j=1}^{5}$ following [23]. Then, the fused feature map $\mathbf{F}_{\mathrm{fuse}}^{i}$ is obtained through the frequency fusion module ($\mathsf{FFM}$), defined as $\mathbf{F}_{\mathrm{fuse}}^{i}=\mathsf{FFM}(\mathsf{concat}(\mathbf{F}_{\mathrm{f}}^{i},\mathsf{ROI}(\mathbf{F}_{\mathrm{c}},\mathbf{P}^{i}))).$ Subsequently, $\mathsf{concat}(\mathbf{F}_{\mathrm{fuse}}^{i},\mathsf{ROI}(\mathbf{F}_{\mathrm{c}},\mathbf{P}^{i}),\mathbf{F}_{\mathrm{enc}})$ is processed through two consecutive layers, each consisting of a $3\times 3$ convolution, batch normalization, and ReLU activation. Finally, the resulting feature is fed into the DPT decoder [31] to obtain the residual map $\mathbf{R}$.

### v3

# One Look is Enough: Seamless Patchwise Refinement for Zero-Shot Monocular Depth Estimation on High-Resolution Images

@@ Abstract
Figure 1: Qualitative comparison of patch-based depth estimation models (PatchFusion [23], PatchRefiner [24] and PRO (Ours)) for high-resolution images. PatchRefiner [24] and PRO (Ours) also visualize the residuals to effectively highlight depth discontinuity issues in the 4th and 6th columns. All models perform depth estimations on $\times 4$ 4 patches of input images. Zoom-in can help distinguish the grid lines (depth discontinuities) of the $\times 4$ 4 patches. Our proposed PRO achieves smooth boundary transitions without depth discontinuity artifacts along the grid boundaries.
@@ 1 Introduction
To address the aforementioned problems, we propose Patch Refine Once (PRO), a novel refinement model that enables DE models to generate accurate depth predictions with only a single refinement per patch at inference time (i.e.,“One Look”). Our PRO consists of two strategies: Grouped Patch Consistency Training and Bias Free Masking. Fig. One Look is Enough: Seamless Patchwise Refinement for Zero-Shot Monocular Depth Estimation on High-Resolution Images shows a qualitative comparison of patch-based depth estimation models for high-resolution images. As shown, our PRO yields depth refinement results with minimal depth discontinuities along grid boundaries, whereas two recent state-of-the-art (SOTA) models still exhibit noticeable boundary artifacts. Fig. 2 illustrates a comparison of performance and efficiency between recent SOTA depth refinement models and our PRO. As observed, our PRO provides the best performance in terms of $\mathrm{D}^{3}\mathrm{R}$, AbsRel and inference speed. Our contribution is summarized as:
@@ 4.2 Implementation Details
We adopt the pretrained DepthAnythingV2 (DA2) [47] as the baseline model ($\Psi$ in Fig. 4). PRO, PatchFusion [23], and PatchRefiner [24] are (re)trained based on DA2, while BoostingDepth [26] uses DA2 without additional training. The input resolution to $\Psi$ is fixed at $\times 518$ 518. When we use the GPCT strategy, adjacent patches are cropped with an overlap of 224 pixels, ensuring an overlap of $h\times 224$ for horizontally adjacent patches and $224\times w$ for vertically adjacent patches, where $h$ and $w$ denote the height and width of the patch, respectively. In our BFM, the dilation kernel size is empirically set to (10, 20). Training is conducted on 7,592 UnrealStereo4K samples for 8 epochs with a batch size of 64 (with gradient accumulation), taking 10 hours on a single RTX 4090 GPU. Additional details are described in supplementary material.
@@ 4.3 Performance Comparison
At test time, each input image is divided into a $\times 4$ 4 grid of patches. The depth of each patch is refined individually, and the refined patches are then reassembled to generate the final depth map.
@@ Appendix A Extra Experiments
$\times 3$ 3 0.0305 0.994 0.0423 0.985 0.0288 0.995 0.0805 0.125
$\times 4$ 4 0.0304 0.994 0.0422 0.985 0.0287 0.996 0.0803 0.156
$\times 5$ 5 0.0305 0.994 0.0423 0.985 0.0288 0.996 0.0797 0.162
@@ Appendix B Architecture of the Fusion Module
The encoder takes $\mathbf{P}^{i}$, $\mathbf{D}_{\mathrm{f}}^{i}$, and $\mathsf{ROI}(\mathbf{D}_{\mathrm{c}},\mathbf{P}^{i})$ as inputs and produces a set of five intermediate features, denoted as $\mathbf{F}_{\mathrm{enc}}=\{\mathbf{f}_{\mathrm{enc},j}^{i}\}_{j=1}^{5}$ following [23]. Then, the fused feature map $\mathbf{F}_{\mathrm{fuse}}^{i}$ is obtained through the frequency fusion module ($\mathsf{FFM}$), defined as $\mathbf{F}_{\mathrm{fuse}}^{i}=\mathsf{FFM}(\mathsf{concat}(\mathbf{F}_{\mathrm{f}}^{i},\mathsf{ROI}(\mathbf{F}_{\mathrm{c}},\mathbf{P}^{i}))).$ Subsequently, $\mathsf{concat}(\mathbf{F}_{\mathrm{fuse}}^{i},\mathsf{ROI}(\mathbf{F}_{\mathrm{c}},\mathbf{P}^{i}),\mathbf{F}_{\mathrm{enc}})$ is processed through two consecutive layers, each consisting of a $\times 3$ 3 convolution, batch normalization, and ReLU activation. Finally, the resulting feature is fed into the DPT decoder [31] to obtain the residual map $\mathbf{R}$.